OpenStudy (anonymous):

Can somebody help me? My teacher didn't help me with this. Will medal for best answer! =^ x ^= y = x^2 + 7x + 100 y +10x = 30

OpenStudy (anonymous):

Substitute (30 - 10x) in the place of y in the first equation.

OpenStudy (anonymous):

20x?

OpenStudy (anonymous):

No! 30 - 10x = x^2 + 7x + 100

OpenStudy (anonymous):

x^2 + 7x +10x + 100 - 30 = 0

OpenStudy (anonymous):

You should be able to solve that.

OpenStudy (nincompoop):

what was being asked tho?

OpenStudy (anonymous):

I'm not quite sure. ._.

OpenStudy (anonymous):

I just had to find both x and y-values and my teacher went out of office

OpenStudy (nincompoop):